Please help me solve this problem. Measures of central tendency. After four algebra tests, Alisa has a mean score of 82. One more 100-point test is to be given in the class. All the test scores are equal importance. Is it possible for Alisa to raise her average(mean) to 90? Please explain. Thank you.
Answer by Theo(13342)

5 test scores with a mean score of 90 would total 450.
she has 4 test scores with a mean of 82 which means the 4 score total is 4 * 82 equals 328.
if she got a hundred on the fifth test, she would total 428 which is less than the 450 required so there's no way she can raise her average score to 90 when there's only 1 test left.
the equation for her last test score would be:
((4 * 82) + x) / 5 = 90
multiply both sides of this equation by 5 to get:
((4*82) + x) = 5 * 90
simplify to get:
328 + x = 450
subtract 238 from both sides of this equation to get:
x = 450 - 328 which simplifies to:
x = 122
she would have to get a 122 on the fifth test to have an average of 90 on all 5 tests.
(4 * 82 + 122) / 5 = 450 / 5 = 90